What number is 16 and two thirds % more than 636?
Answer #1

Answer:

16 and two thirds % = (16 + 2/3)% = ((48 + 2) /3)% = (50/3)%

The number that is 16 and two thirds % more than 636 is = 636 * (1 + (50/3)%) = 742

The number that is 16 and two thirds % more than 636 is = 742